# Read Anywhere Pointed: Layout-aware GUI Screen Reading with Tree-of-Lens Grounding

### v1

@@
Graphical User Interfaces (GUIs) are central to our interaction with digital devices. Recently, growing efforts have been made to build models for various GUI understanding tasks. However, these efforts largely overlook an important GUI-referring task: screen reading based on user-indicated points, which we name the Screen Point-and-Read (ScreenPR) task. This task is predominantly handled by rigid accessible screen reading tools, in great need of new models driven by advancements in Multimodal Large Language Models (MLLMs). In this paper, we propose a Tree-of-Lens (ToL) agent, utilizing a novel ToL grounding mechanism, to address the ScreenPR task. Based on the input point coordinate and the corresponding GUI screenshot, our ToL agent constructs a Hierarchical Layout Tree. Based on the tree, our ToL agent not only comprehends the content of the indicated area but also articulates the layout and spatial relationships between elements. Such layout information is crucial for accurately interpreting information on the screen, distinguishing our ToL agent from other screen reading tools. We also thoroughly evaluate the ToL agent against other baselines on a newly proposed ScreenPR benchmark, which includes GUIs from mobile, web, and operating systems. Last but not least, we test the ToL agent on mobile GUI navigation tasks, demonstrating its utility in identifying incorrect actions along the path of agent execution trajectories. Code and data: screen-point-and-read.github.io.
@@ 1 Introduction
Graphical User Interfaces (GUIs), dominate our digital interactions with visually rich screenshots featuring colors, icons, texts, and spatial layouts. Recognizing that screenshots are more accessible and intuitive for depicting visual cues, there are growing efforts to build AI agents to interpret GUIs visually Shaw et al. (2023); Deng et al. (2023); Cheng et al. (2024); Hong et al. (2023); You et al. (2024). Existing works, however, overlook the point-based screen reading task, where the input is the coordinate of a user’s indicated point on the screen and the corresponding screenshot, and the output is a descriptive interpretation of the screen region pointed. This task is critical for accessible technology, providing valuable assistance to users with visual impairments and we refer to such a task as the Screen Point-and-Read (ScreenPR) task.
Aiming to solve the ScreenPR task, we introduce the Tree-of-Lens (ToL) agent. Taking advantage of the generalizability of the advanced Multimodal Large Language Models (MLLMs), the ToL agent is built for GUIs screenshots across domains with a point anywhere on the screen as input. The output from the ToL agent is natural language descriptions for both the content of the indicated region and the related information about the screen layout, as shown in Figure 1. Especially, we find that it is critical to articulate the layout between elements in the screenshot for users to gain a comprehensive understanding of the interface and avoid ambiguity. For example, with only content description, one cannot distinguish the two identical “Tumbler pack" shown in Figure 1.
@@
In addition, we also test our ToL agent with trajectories from a mobile GUI navigation agent. By applying ToL agent to describe the point of each action taken in the trajectory, we demonstrate its utility in identifying incorrect actions along the execution path. This capability is pivotal for refining the development of mobile agents, as it provides clear, actionable feedback on navigation decisions. This application not only validates the ToL agent’s effectiveness but also opens up possibilities for strengthening the performance of mobile agents with sophisticated, layout-aware GUI understanding capabilities.
@@ 2 Related Works
Recent advancements in Multimodal Large Language Models (MLLMs) have led to arising a series of highly capable generalist models Liu et al. (2024b); Chen et al. (2023); Ye et al. (2023); Dai et al. (2023); Bai et al. (2023); Liu et al. (2024a). Spurred by the strong general capability of MLLMs, recent works have been focused on utilizing and improving MLLMs for GUI tasks. Some focus mainly on a specific GUI task such as GUI navigation Wang et al. (2024); Yang et al. (2023b); Zheng et al. (2024), GUI referring expression comprehension Cheng et al. (2024), while some aim at a set of GUI grounding and referring tasks simultaneously, such as CogAgent Hong et al. (2023), which is Multi-task Fine-tuned on CogVLM Wang et al. (2023), and Ferret-UI You et al. (2024), ensuing Ferret You et al. (2023); Zhang et al. (2024), tailored for GUI domain. In our work, we focus on the Screen Point-and-Read (ScreenPR) task, a referring task in which inputs are point locations on the screen with the corresponding screenshots, and develop a specific benchmark to evaluate the model’s performance in such tasks. Also, inspired by big challenges that MLLMs are facing in understanding images with complex layout information Fan et al. (2024), we propose the Tree-of-Lens (ToL) agent that generates layout-aware description, which distinguishes from all other GUI agents with GUI referring ability.
@@ 3.3 Hierarchical Layout Tree Construction
3.4 Target Path Selection and Multi-lens Prompting
@@
Then a Multi-lens Prompting method is adopted, where the curated set of lenses is sent to GPT-4o with a text prompt which includes the text coordinate of the target point. Some specific instructions are also included in the prompt, asking, "Explain where box 1 is in relation to box 2 and where box 2 is located on the overall screen." In this way, we provide clear visual cues that facilitate the GPT-4o to generate detailed descriptions that reflect the GUI’s content and layout. For more details about the prompt used, please refer to Appendix C.
@@ 4.1 Overview
We introduce the Screen Point-and-Read (ScreenPR) benchmark to rigorously evaluate our ToL agent on the ScreenPR task, where descriptions are required based on user-indicated points on screenshots. As shown in Table 1, this benchmark covers a diverse domain of GUIs with a total 650 screenshots. Each screenshot is annotated with about 2 target points and their corresponding local regions on the screenshot. Figure 4i illustrates the normalized coordinates of target points $P_{i}$ in the ScreenPR benchmark, demonstrating thorough and comprehensive coverage of every location on the screen. The regions included have various content, and their area ranges from 1% to over 50% of the screen, as shown in Figure 4ii.
@@ 4.2 Data Collection
We first gather mobile and web screenshots from Screenspot Cheng et al. (2024) and operating system screenshot from the data explorer videos [^3] https://os-world.github.io/explorer.html of OSWorld Xie et al. (2024). Then we randomly generate 3 or 4 candidates of the target point $P_{i}$ for each screenshot. Then, to maximize the variety of the region indicated by points, we hire 3 students to check these candidate points, where they assess all the candidate points on each screenshot and remove those that represent duplicated regions. Then, the remaining points are regarded as the target point $P_{i}$, and the annotators draw the bounding boxes of the corresponding local regions $R_{i}$. After the annotation, scripts were used to spot errors, such as ensuring that local regions $R_{i}$ included $P_{i}$ inside. The annotated data is then manually examined by authors to ensure quality.
@@ 4.3 Evaluation
We conduct surveys with human judges to assess the quality of the generated descriptions. Each survey presents a screenshot with an indicated point, accompanied by descriptions from different screen readers. The judges are asked to rate these descriptions based on three criteria: how well they describe the content, how well they articulate the layout, and their overall preference. Each question had four options, ranging from "very well" to "not at all," which we numerically mapped to accuracy scores of 100%, 66%, 33%, and 0%, respectively.
@@
In order to evaluate screen reader models automatically, we propose the cycle consistency evaluation. The outputs $(\hat{D^{c}},\hat{D^{l}})$ from the screen-reader are each used as the input for an auxiliary model tasked with specific validation tasks, with the alternative model’s performance indicating the quality of the screen-reader’s descriptions.
Specifically, to evaluate content accuracy, GPT-4o serves as the auxiliary model, performing a multi-choice selection task. This alternative model is asked to select one from the multi-choice candidates based on the predicted descriptions $\hat{D^{c}}$. The candidates are four screenshots of different GUI regions cropped based on human annotation, with only one targeted region
@@
For layout description accuracy, we design another multi-choice selection task for the auxiliary model, GPT-4-turbo. This multi-choice selection task requires the auxiliary model to select the positional relationship of two regions indicated by $P_{i}$ and $P_{i^{\prime}})$ based on corresponding generated layout descriptions $\hat{D^{l}}_{i}$ and $\hat{D^{l}}_{i^{\prime}}$. $P_{i^{\prime}}$, the reference point, is manually annotated to represent a region $R_{i^{\prime}}$ that does not overlap with the local region of the target point $R_{i}$ on the same screen $S_{i}$, especially for the evaluation of the layout description accuracy. The multi-choice selection task presented to the auxiliary model is
@@
Table 2: Main results. We evaluate our ToL agent against three other baselines with the human evaluation and cycle consistency evaluation of our ScreenPR benchmark. Additionally, we compare the generated content descriptions with human verified content descriptions using language similarity scores. The results show that the ToL agent achieves the best performance.
@@
+ w/o target point mark 65.6  (-9.4) 37.6 (-2.1)
+ w/o local & global regions marks 35.1 (-29.9) 21.9 (-18.6)
Table 3: Ablation results. We gradually refine the ablation to exclude various processes in our ToL grounding mechanism for ToL agent on the ScreenPR benchmark. The results show the effectiveness of various components.
@@ 5 Experiments
We adopt several baselines, including generalist MLLMs: the proprietary GPT-4o OpenAI (2024) and the open-source LlaVA-NeXT Liu et al. (2024a). We also test the CogAgent Hong et al. (2023), a MLLM specifically tuned for GUI tasks. The inputs for baselines are the text prompt with the input point coordinate $P_{i}$ and the screenshot $S_{i}$. We keep prompts for baselines in the same style as the ToL agent, expect that we have to simplify it for CogAgent as we find it necessary for CogAgent to follow the instructions. Detailed prompt design is shown in Appendix C.
@@ 5.1 Main Results
Additionally, we leverage language similarity scores, including BERTScore and ROUGE-L, to compare the generated content descriptions with human-verified descriptions. These descriptions are captions generated using GPT-4V with the direct input of $Crop(S_{i},R_{i})$, and we manually examined and corrected mistakes. We observe that the overall trend from the language similarity evaluation results matches the content evaluation results from human and cycle consistency evaluations. However, the differences in language similarity scores between models are relatively indistinguishable. We believe this is because the captions are more comprehensive, sometimes redundant, whereas the screen reading output is more focused on delivering the key features of the content to help users understand it.
@@ 5.3 Broader Application: Verification of Mobile Navigation Agent Actions
In this section, we demonstrate how our ToL agent helps identify incorrect actions for a mobile navigation agent, MagicWonder, on MagicWand platformDing et al. . Driven by a specific task goal, such as "connect to Wifi in settings", MagicWonder can generate actions to interact with a live mobile screen and save the detailed information for each step along its execution trajectory. More details are provided in Appendix D.1. To incoorporate our ToL agent with the verification purpose, we design a simple pipeline. As shown in Figure 5, the layout and content descriptions are generated by our ToL agent for each action point and sent to the GPT-4-Turbo model. The GPT-4-Turbo model, based on the task goal, instructions, and descriptions from ToL agent, decides whether the last action from the MagicWonder is correct or not.
@@
We select 52 mobile agent trajectories with specific task goals from instruction-guided executions on MagicWand platformDing et al. , with more details in Appendix D.2. Each trajectory consists of multiple execution steps triggered by distinct actions, resulting in a total action number of 209. We hire two students to annotate incorrect actions in the chosen trajectories to get the ground truth labels.
@@
We perform 10 end-to-end evaluations and present the mean value and standard deviation in Table 5. Despite minor fluctuations, while 92% correct actions can be stably recognized, about 38% incorrect actions can be identified. This reveals that incorporating the description from the ToL agent can significantly and robustly improve mobile navigation agents’ capability. Our ToL agent can also identify particularly tricky challenges faced by mobile navigation agents during execution. One notable issue is the "execution loop," where agents repeatedly take actions on the same or similar regions without making progress. Our results show that for all 20 incorrect actions fall into this category, 44% ± 0.04 can be successfully detected.
@@
6 Conclusion and Discussion
In this work, we propose a novel Tree-of-Lens (ToL) grounding method and build a ToL agent for the Screen Point-and-Read (ScreenPR) task, an important screen referring task. Our ToL agent outputs layout-aware descriptions for pointed regions on screenshots. With our newly proposed Screen Point-and-Read (ScreenPR) benchmark, we show that our ToL agent outperforms all other baselines in terms of content and layout description accuracy, demonstrating great potential to improve accessibility of digital devices for visually impaired users. We also apply the ToL agent to identify incorrect actions in the execution trajectories of a mobile navigation agent, achieving promising results.

### v2

@@
Graphical User Interfaces (GUIs) are central to our interaction with digital devices and growing efforts have been made to build models for various GUI understanding tasks. However, these efforts largely overlook an important GUI-referring task: screen reading based on user-indicated points, which we name the Screen Point-and-Read (ScreenPR) task. Currently, this task is predominantly handled by rigid accessible screen reading tools, in great need of new models driven by advancements in Multimodal Large Language Models (MLLMs). In this paper, we propose a Tree-of-Lens (ToL) agent, utilizing a novel ToL grounding mechanism, to address the ScreenPR task. Based on the input point coordinate and the corresponding GUI screenshot, our ToL agent constructs a Hierarchical Layout Tree. Based on the tree, our ToL agent not only comprehends the content of the indicated area but also articulates the layout and spatial relationships between elements. Such layout information is crucial for accurately interpreting information on the screen, distinguishing our ToL agent from other screen reading tools. We also thoroughly evaluate the ToL agent against other baselines on a newly proposed ScreenPR benchmark, which includes GUIs from mobile, web, and operating systems. Last but not least, we test the ToL agent on mobile GUI navigation tasks, demonstrating its utility in identifying incorrect actions along the path of agent execution trajectories. Code and data: https://screen-point-and-read.github.io.
@@ 1 Introduction
Graphical User Interfaces (GUIs), dominate our digital interactions with visually rich screenshots featuring colors, icons, texts, and spatial layouts. Recognizing that screenshots are more accessible and intuitive for depicting visual cues, there are growing efforts to build AI agents to interpret GUIs visually Shaw et al. (2023); Deng et al. (2023); Cheng et al. (2024); Hong et al. (2023); You et al. (2024). Existing works, however, overlook the point-based screen reading task, where the input is the coordinate of a user’s indicated point on the screen with the corresponding screenshot, and the output is a descriptive interpretation of the screen region pointed. This task is critical for accessible technology, providing valuable assistance to users with visual impairments and we refer to such a task as the Screen Point-and-Read (ScreenPR) task.
Aiming to solve the ScreenPR task, we introduce the Tree-of-Lens (ToL) agent. Taking advantage of the generalizability of the advanced Multimodal Large Language Models (MLLMs), the ToL agent is built for GUI screenshots across domains with a point anywhere on the screen as input. The output from the ToL agent is natural language descriptions for both the content of the indicated region and the related information about the screen layout, as shown in Figure 1. Especially, we find that it is critical to articulate the layout between elements in the screenshot for users to gain a comprehensive understanding of the interface and to avoid ambiguity. For example, with only content description, one cannot distinguish the two identical “Tumbler pack" shown in Figure 1.
@@
In addition, we also test our ToL agent with trajectories from a mobile GUI navigation agent. By applying the ToL agent to describe the point of each action taken in the trajectory, we demonstrate its utility in identifying incorrect actions along the execution path. This capability is pivotal for refining the development of mobile agents, as it provides clear, actionable feedback on navigation decisions. This application not only validates the ToL agent’s effectiveness but also opens up possibilities for strengthening the performance of mobile agents with sophisticated, layout-aware GUI understanding capabilities.
@@ 2 Related Works
Recent advancements in Multimodal Large Language Models (MLLMs) have led to a rising series of highly capable generalist models Liu et al. (2024b); Chen et al. (2023); Ye et al. (2023); Dai et al. (2023); Bai et al. (2023); Liu et al. (2024a). Spurred by the strong general capability of MLLMs, recent works have been focused on utilizing and improving MLLMs for GUI tasks. Some focus mainly on a specific GUI task such as GUI navigation Wang et al. (2024); Yang et al. (2023b); Zheng et al. (2024), GUI referring expression comprehension Cheng et al. (2024), while some aim at a set of GUI grounding and referring tasks simultaneously, such as CogAgent Hong et al. (2023), which is Multi-task Fine-tuned on CogVLM Wang et al. (2023), and Ferret-UI You et al. (2024), ensuing Ferret You et al. (2023); Zhang et al. (2024), tailored for GUI domain. In our work, we focus on the Screen Point-and-Read (ScreenPR) task, a referring task in which inputs are point locations on the screen with the corresponding screenshots, and develop a specific benchmark to evaluate the model’s performance in such tasks. Also, inspired by big challenges that MLLMs are facing in understanding images with complex layout information Fan et al. (2024), we propose the Tree-of-Lens (ToL) agent that generates layout-aware description, which distinguishes from all other GUI agents with GUI referring ability.
@@ 3.3 Hierarchical Layout Tree Construction
3.4 Target Path Selection and Multi-lense Prompting
@@
Then a Multi-lens Prompting method is adopted, where the curated set of lenses is sent to GPT-4o with a text prompt which includes the text coordinate of the target point. Some specific instructions are also included in the prompt, asking, "Explain where box 1 is in relation to box 2 and where box 2 is located on the overall screen." In this way, we provide clear visual cues that facilitate the GPT-4o to generate detailed descriptions that reflect the GUI’s content and layout. For more details about the prompt used, please refer to Appendix D.
@@ 4.1 Overview
We introduce the Screen Point-and-Read (ScreenPR) benchmark to rigorously evaluate our ToL agent on the ScreenPR task, where descriptions are required based on user-indicated points on screenshots. As shown in Table 1, this benchmark covers a diverse domain of GUIs with a total 650 screenshots. Each screenshot is accompanied by manually verified target points $P_{i}$, their corresponding local regions $R_{i}$, and content descriptions. Figure 4i illustrates the normalized coordinates of target points $P_{i}$ in the ScreenPR benchmark, demonstrating comprehensive coverage of every location on the screen. The regions included have various content, and their area ranges from 1% to over 50% of the screen, as shown in Figure 4ii.
@@ 4.2 Data Collection
We first gather mobile and web screenshots from Screenspot Cheng et al. (2024) and operating system screenshots from the data explorer videos [^3] https://os-world.github.io/explorer.html of OSWorld Xie et al. (2024). Then we randomly generate 4 or 5 points as the candidates for target points $P_{i}$ within each screenshot. Next, we hire 3 students as annotators to draw bounding boxes for the local regions $R_{i}$ corresponding to the points. To maximize the variety of the data, candidate target points that represent duplicated local regions will be removed. Finally, we assign the remaining points within each screenshot as either target points $P_{i}$ or reference points $P_{i^{\prime}}$, where they are paired such that every pair of $(P_{i},P_{i^{\prime}})$ all have non-overlapping corresponding local regions ($P_{i^{\prime}}$ is used for our proposed cycle consistency evaluation detailed in Section 4.3). Additionally, we collect human-verified descriptions for all local regions $R_{i}$ corresponding to $P_{i}$. These descriptions are captions generated using GPT-4o with the direct input of $Crop(S_{i},R_{i})$, and we manually examined and corrected mistakes. After the annotation, we ensure the data quality by first using scripts to spot errors, such as ensuring that local regions $R_{i}$ are included $P_{i}$ inside, and then manually examine the data.
@@ 4.3 Evaluation
We conduct surveys with human judges to assess the quality of the generated descriptions. Each survey presents a screenshot with an indicated point, accompanied by descriptions from different screen readers. The judges are asked to rate these descriptions based on three criteria: how well they describe the content, how well they articulate the layout, and their overall preference. Each question has four options, ranging from "very well" to "not at all," which we numerically mapped to accuracy scores of 100%, 66%, 33%, and 0%, respectively.
@@
In order to evaluate screen reader models automatically, we propose the cycle consistency evaluation. The outputs $(\hat{D^{c}},\hat{D^{l}})$ from the screen reader are each used as the input for an auxiliary model tasked with specific validation tasks, with the alternative model’s performance indicating the quality of the screen reader’s descriptions.
Specifically, to evaluate content accuracy, GPT-4V Achiam et al. (2023) serves as the auxiliary model, performing a multi-choice selection task. This alternative model is asked to select one from the multi-choice candidates based on the predicted descriptions $\hat{D^{c}}$. The candidates are four screenshots of different GUI regions cropped based on human annotation, with only one targeted region
@@
For layout description accuracy, we design another multi-choice selection task for the auxiliary model, GPT-4 Achiam et al. (2023). This multi-choice selection task requires the auxiliary model to select the positional relationship of two regions indicated by $P_{i}$ and $P_{i^{\prime}}$ based on corresponding generated layout descriptions $\hat{D^{l}}_{i}$ and $\hat{D^{l}}_{i^{\prime}}$. $P_{i^{\prime}}$, the reference points, collected in the same way as target points $P_{i}$ (detailed in Section 4.2) are manually annotated to represent a region $R_{i^{\prime}}$ that does not overlap with the local region of the target point $R_{i}$ on the same screen $S_{i}$, especially for the evaluation of the layout description accuracy. The multi-choice selection task presented to the auxiliary model is
@@
Table 2: Main results. We evaluate our ToL agent against three other baselines with the human evaluation and cycle consistency evaluation of our ScreenPR benchmark. Additionally, we compare the generated content descriptions with human-verified content descriptions using language similarity scores. The results show that the ToL agent achieves the best performance.
Content Acc Layout Acc
GPT-4o 35.10 21.87
ToL Agent (GPT-4o) 74.96 39.67
Qwen-VL-Plus 28.29 12.50
ToL Agent (Qwen-VL-Plus) 41.93 14.47
Gemini Pro Vision 33.20 12.27
ToL Agent (Gemini Pro Vision) 45.33 18.53
Table 3: Results of the ToL Agent built upon various MLLMs. The results of our ToL agent based on various MLLMs are consistently better than the MLLMs alone, providing strong evidence that our approach is effective and generalizable. All the MLLMs are accessed through their APIs around Jun 2024.
@@
+ w/o target point mark ($\approx$ SoM) 65.6  ($-9.4$) 37.6 ($-2.1$)
 + w/o local & global regions marks
(degrade to vanilla GPT-4o)
35.1 ($-29.9$) 21.9 ($-18.6$)
Table 4: Ablation results. We gradually refine the ablation to exclude various processes in our ToL grounding mechanism for the ToL agent on the ScreenPR benchmark. The results show the effectiveness of various components.
@@ 5 Experiments
We adopt several baselines, including generalist MLLMs: the proprietary GPT-4o OpenAI (2024) and the open-source LlaVA-NeXT Liu et al. (2024a). We also test the CogAgent Hong et al. (2023), a MLLM specifically tuned for GUI tasks, and a visual prompting method Scaffolding Lei et al. (2024) that is specially developed to enhance the spacial reasoning ability of MLLMs. The inputs for baselines are the text prompt with the input point coordinate $P_{i}$ and the screenshot $S_{i}$. We keep prompts for baselines in the same style as the ToL agent, except that we have to simplify it for CogAgent as we find it necessary for CogAgent to follow the instructions. Detailed prompt design is shown in Appendix D.
@@ 5.1 Main Results
Additionally, we leverage language similarity scores, including BERTScore and ROUGE-L, to compare the generated content descriptions with human-verified descriptions. We observe that the overall trend from the language similarity evaluation results matches the content evaluation results from human and cycle consistency evaluations. However, the differences in language similarity scores between models are relatively indistinguishable. We believe this is because the captions are more comprehensive, sometimes redundant, whereas the screen reading output is more focused on delivering the key features of the content to help users understand it.
To further evaluate the generalizability of our ToL agent pipeline, we conduct experiments using the cycle consistency evaluation on our ToL agent with two other different MLLMs rather than the default GPT-4o: the Qwen-VL-Plus model Bai et al. (2023) and the Gemini Pro Vision model Team et al. (2023). We access both of these models via their APIs. As shown in Table 3, the results show that our ToL agent built upon various MLLMs consistently outperforms the standalone MLLMs, offering compelling evidence that our approach is effective and generalizable.
@@ 5.3 Broader Application: Verification of Mobile Navigation Agent Actions
In this section, we demonstrate how our ToL agent helps identify incorrect actions for a mobile navigation agent, MagicWonder, on the MagicWand platform Ding et al. . Driven by a specific task goal, such as "connect to Wifi in settings", MagicWonder can generate actions to interact with a live mobile screen and save the detailed information for each step along its execution trajectory. More details are provided in Appendix E.1. To incorporate our ToL agent for verification of MagicWonder’s output actions, we design a simple pipeline. As shown in Figure 5, the layout and content descriptions are generated by our ToL agent for each action-point-screenshot triplet and sent to the GPT-4 model. The GPT-4 model, based on the task goal, instructions, and descriptions from the ToL agent, decides whether the last action from the MagicWonder is correct.
@@
We select 52 mobile agent trajectories with specific task goals from instruction-guided executions on the MagicWand platform, with more details in Appendix E.2. Each trajectory consists of multiple execution steps triggered by distinct actions, resulting in a total action number of 209. We hire two students to annotate incorrect actions in the chosen trajectories to get the ground truth labels.
@@
We perform 10 end-to-end evaluations and present the mean value and standard deviation in Table 6. Despite minor fluctuations, while 92% correct actions can be stably recognized, about 38% incorrect actions can be identified. This reveals that by incorporating the ToL agent, the performance of the mobile navigation agent is significantly and robustly improved. Our ToL agent can also identify particularly tricky challenges faced by mobile navigation agents during execution. One notable issue is the "execution loop", where agents repeatedly take actions on the same or similar regions without making progress. Our results show that for all 20 incorrect actions that fall into this category, 44% ± 0.04 of them can be successfully detected. To further validate the effectiveness of our method, we introduce two baselines: Confidence Score Filter and GPT-4o Verifier. The Confidence Score Filter considers actions as positive when the relevant confidence score is above 0.7 and the GPT-4o Verifier directly sends the actions with corresponding screenshots as prompts to GPT-4o to identify the correctness of actions. We run the end-to-end evaluation equivalently and the results in Table 7 show that our method of using the ToL agent to verify the output from the mobile navigation agent outperforms the other baselines.
@@
Verification Method F1 score Repetition Detection Rate
Confidence Score Filter 65%$\pm$0.00 0%$\pm$0.00
GPT-4o 61%$\pm$0.01 40%$\pm$0.10
ToL Agent 74%$\pm$0.01 44%$\pm$0.04
Table 7: Performance of different verification methods in verifying actions of mobile navigation agent. The result shows that using our ToL agent to verify the actions of the mobile navigation agent achieves the best F1 score and performs the best in detecting the notable issue of the "execution loop", i.e. repetition detection.
6 Conclusion
In this work, we propose a novel Tree-of-Lens (ToL) grounding method and build a ToL agent for the Screen Point-and-Read (ScreenPR) task, an important GUI referring task that is critical for accessible technology. Our ToL agent outputs layout-aware descriptions for pointed regions on screenshots. With our newly proposed Screen Point-and-Read (ScreenPR) benchmark, we show that our ToL agent outperforms all other baselines in terms of content and layout description accuracy, demonstrating great potential in assisting visually impaired users when using digital devices. Last but not least, in our experiment, we also apply our ToL agent to identify incorrect actions performed by a mobile navigation agent, yielding promising results.